# WebDRO: A Web-based Group-level Clustering and Reweighting Method for Unsupervised Dense Retrieval

### v3

@@ Abstract
The anchor-document data derived from web graphs offers a wealth of paired information for training dense retrieval models in an unsupervised manner. However, the presence of inherent noise invariably compromises the robustness of training dense retrieval models, consequently hurting the performance. In this paper, we introduce WebDRO, an efficient approach for clustering the web graph data and optimizing group weights to enhance the robustness of the pretraining process of dense retrieval models on web graphs. Initially, we build an embedding model for clustering anchor-document pairs. Specifically, we contrastively train the embedding model for link prediction, which guides the embedding model in capturing the inherent document features behind the web graph links. Subsequently, we employ the group distributional robust optimization to recalibrate the weights across different clusters of anchor-document pairs during training dense retrieval models, directing the model to assign higher weights to clusters with higher loss and focus more on worst-case scenarios. Our experiments conducted on MS MARCO and BEIR demonstrate that our method can effectively improve retrieval performance in unsupervised training settings. Further analysis confirms the stability and validity of group weights learned by WebDRO. All codes will be released via GitHub.
WebDRO: A Web-based Group-level Clustering and Reweighting Method for Unsupervised Dense Retrieval
@@ 1 Introduction
Dense retrieval models, which encode the semantics of text into dense vectors, have become prevalent and achieved state-of-the-art performance on many public benchmarks thanks to the surge of pre-trained language models like BERT Devlin et al. (2019) and T5 Raffel et al. (2020). With enough annotated data, dense retrieval models can adeptly capture semantic features in supervised scenarios, outperforming traditional sparse methods such as BM25 Robertson et al. (2009, 1995), which solely retrieves documents containing keywords already present in the query. However, annotated data is inaccessible in many scenarios either due to the extensive effort needed for annotation or privacy concerns. This highlights the necessity for unsupervised dense retrievers. Several paradigms of unsupervised dense retrieval have been proposed, among which leveraging web anchors is a promising choice. Specifically, anchor-document pairs can approximate supervised signals where the document contains relevant information that appears in the query better than most existing methods Xie et al. (2023a); Zhang et al. (2020b).
Web anchors contain both useful signals and noise, limiting the effectiveness of anchor-based training methods Zhang et al. (2020b). Previous research has primarily focused on instance-level selection to distinguish useful contrastive pairs from uninformative ones by assigning a weight to each instance Moore and Lewis (2010); Zhang et al. (2020a). Nevertheless, directly weighting at the instance level is often coarse, since the information contained in individual data points is insufficient. To reach a more tailored weighting, some approaches leverage feedback signals during the training process, such as Active Learning Settles (2011); Coleman et al. (2020) and Meta Adaptive Ranking Sun et al. (2021). However, these methods have the drawback of requiring the model to be trained multiple times, which is highly time-consuming and computationally intensive.
@@
We address the lack of both efficient and accurate retrieval data selection methods by introducing a group-level clustering and reweighting strategy–WebDRO. WebDRO enables us to conduct data selection with the feedback signal (training loss) in a single pass, as well as, cluster data into groups due to that the group encapsulates significantly more information than a single data point. In other words, we are “smoothing” the weights within each group through clustering, thereby enhancing the quality and stability of reweighting. Specifically, we employ a two-step approach to train an unsupervised dense retrieval model using WebDRO. In the first step, we utilize structural information from the web graph to train an embedding model with contrastive learning, which plays a role as the link predictor. By leveraging the embedding model to conduct vector representations for documents in the anchor-document pairs, we can produce high-quality clusters of contrastive pairs. In the second step, we utilize Group Distributionally Robust Optimization (GroupDRO) Oren et al. (2019); Sagawa et al. (2019) to reweight the clusters. Originally designed to tackle distributional shifts in language model training, GroupDRO dynamically adjusts the weights of different groups based on the difficulty of learning with that group. Consequently, the model can effectively discern informative web pages from noisy, uninformative ones.
@@
We conduct experiments on two widely used retrieval datasets, MS MARCO Nguyen et al. (2016) and BEIR Thakur et al. (2021). Results indicate that our method notably enhances retrievers’ performance in both unsupervised and finetuned settings. Specifically, WebDRO yields a performance improvement of 1.2% (measured by nDCG@10) on MS MARCO and 1.1% on BEIR compared to our main baseline, Anchor-DR Xie et al. (2023b), in an unsupervised setting. Analytical experiments further demonstrate that our model consistently produces stable, high-quality clusters and accurately identifies groups of high importance. Additionally, our method exhibits notable efficiency and scalability advantages, rendering it particularly well-suited for pretraining based on large-scale corpus.
@@ 2 Related Work
This section describes relevant work, focusing on dense retrieval models, data selection methods, and distributionally robust optimization.
Dense Retrieval. Dense Retrieval is an approach where both the query and document are separately encoded through a neural network. The prevalent training approach, introduced by Karpukhin et al. (2020), trains the network with a contrastive loss Although dense retrieval has over-performed sparse methods Robertson et al. (2009) in many tasks, its benefits are generally confined to well-annotated data. Most unsupervised dense retrieval methods focus on using unannotated data to generate query-document pairs that semantically resemble contrastive pairs in supervised scenarios, like ICT Izacard et al. (2021) and COCO-LM Meng et al. (2021). Hyperlinks on the Internet contain rich information about relationships between texts, and thus can also be utilized in dense retrieval. Anchor-DR Xie et al. (2023b) contrastively trains dense retrievers using web anchors-document pairs, which has shown superiority over other unsupervised dense retrieval methods.
Retrieval Data Selection. Most previous work selects retrieval data by comparing the query’s similarity with texts that are considered high-quality. Typically, they assign an attribute to each instance and select a subset based on the attribute’s value Moore and Lewis (2010); Zhang et al. (2020a). However, being similar in form to high-quality text alone does not necessarily make a good query Zhang et al. (2020b). To tailor the selection, some approaches use feedback signals from training as the criterion for selection, namely, Meta Adaptive Ranking Sun et al. (2021) and Active Learning Coleman et al. (2020). The main drawback of using feedback signals is time and computational power consumption since this type of method needs to train the model several times. Such methods are not suitable for pretraining dense retrievers using large-scale datasets.
@@
Figure 1: Training process of WebDRO. Figure 0(a) shows the clustering process. We use web links to train an embedding model and use the model to cluster documents. Figure 0(b) shows the training process of the retrieval model. We aim to align the anchor with the corresponding document while pushing apart unrelated ones. Each group is assigned a weight that is dynamically updated.
@@ 3 Methodology
In this section, we first describe the preliminary of dense retrieval (Sec. 3.1). Then we introduce our WebDRO method, which is shown in Figure 1. To begin with, we use the method in Sec. 3.2 to train an embedding model using the web links. Then, we cluster the anchor-document pairs based on vector representations produced by the embedding model. Finally, we train a robust dense retrieval model with Group Distributionally Robust optimization (GroupDRO) as introduced in Sec. 3.3.
@@ 3.1 Preliminary of Dense Retrieval
In dense retrieval, we train a model to encode both queries and documents as dense vectors and then conduct relevance modeling in an embedding space. Through contrastive learning Oord et al. (2018); Chen et al. (2020); He et al. (2020), the encoder is optimized to align the queries with related documents and pull away the irrelevant documents from queries. We use cosine similarity to measure the relevance between queries and documents and then we minimize the following loss function:
$l_{\text{dr}}(q)=-\log\frac{e^{cos(q,p^{+})/\tau}}{e^{cos(q,p^{+})/\tau}+\sum_{i=1}^{k}{e^{cos(q,p^{-}_{i})/\tau}}},$
@@
3.2 Clustering with the Embedding Model
@@
This subsection introduces our web graph-based anchor-document cluster method. It learns the document representations using the document links in the web graph and then clusters the documents according to their semantic features.
@@
Document Relations in Web Graph. The web graph models the document relationship using hyperlinks, which allows users to navigate from one web page to another or within the same web page. It is evident that the documents linked with a hyperlink will likely share similar or related information. As shown in Figure 0(a), these four documents are related and are linked with the anchor hyperlink, e.g. the link between document A and document B indicate that “horses and gorillas have some similarities”, showing an analogy relationship; the links between document B and document C indicate the definitions of “horses” and “herbivorous”, showing the hypernymy relationship.
@@
Embedding Model Training. We train an embedding model using the document relations in the web graph. We encode the documents using pretrained language models and contrastively train the encoder, which is the same as Sec. 3.1. A web page can link to more than one page, so we enumerate each link to construct contrastive pairs as training examples. For each link, we regard the source web page as the query, the target web page as the positive document, and the unlinked documents as negative ones. As a result, the embedding model is trained as a link predictor, which can distinguish linked web pages from non-linked ones.
@@
Document Clustering. We use the embedding model to encode all documents as dense vectors, and then use Mini-Batch K-Means Sculley (2010), a variant of K-Means Krishna and Murty (1999) to cluster embedding vectors of documents into $n$ groups. Some documents have more than one anchor that links to it, these anchor-document pairs are naturally clustered into the same group. To make sure each group has statistical significance, we set a threshold of the number of documents of different groups as $MinSize=128$. All the groups whose size is smaller than $MinSize$ will be aggregated together and considered together as a larger group. As this large group contains heterogeneous elements with different topics, we should not assign a single weight to this group. Therefore, the large group will not be considered in the reweighting, which is discussed in Sec. 3.3.
3.3 WebDRO: Robustly Unsupervised Retrieval Model Training
The anchor-document pairs contain different types of documents. The noises in these pairs Zhang et al. (2020b) may affect the performance of dense retrievers and conduct unstable training process. To alleviate this problem, we follow Sagawa et al. (2019) to reweight different anchor-document groups and robustly train dense retrievers.
@@ 3.3 WebDRO: Robustly Unsupervised Retrieval Model Training
Following GroupDRO, each group is assigned with a weight that indicates its importance, and the weights are dynamically updated during the training process. Specifically, we divide the anchor-document pairs into $n$ groups after clustering. Then, for a training instance $p$ that belongs to group $k$, the training loss of WebDRO is calculated as:
$L_{\text{WebDRO}}(p)=l_{\text{dr}}(p)\times w_{k}\times n\times C_{k},$ (2)
where $w_{k}$ is the weight of the $k$-th group, $C_{k}$ is size factor of the $k$-th group, and $l_{\text{dr}}$ is the contrastive loss calculated using Eq. 1. In Figure 0(b), we illustrate 3 clusters, each of them containing documents about the same topic. During contrastive training, group weights, which are multiplied onto the training loss, get dynamically updated.
Then, we describe $C_{k}$ and $w_{k}$, respectively.
@@
Size Factor. In a standard GroupDRO scenario, data are sampled from each group according to the uniform distribution. However, the size of different anchor-document groups in our method may vary significantly. To accommodate this circumstance, WebDRO introduces a size factor $C_{k}$:
$C_{k}=\frac{\sum_{j=1}^{n}{N_{j}}}{n\times N_{k}},$ (3)
@@
where $N_{k}$ is the size of the $k$-th anchor-document group. Smaller groups will have a larger $C_{k}$ value, which serves as a balance for different group sizes.
@@
Group Weighting. We denote the weight of group $k$ at step $t$ as $w_{k}^{t}$. The initial weight of all groups is $w_{k}^{0}=\frac{1}{n}$ and after each step, the sum of group weights should be $1$. In each step, the groups’ weights are updated before the calculation of $L_{\text{WebDRO}}$ (Eq. 2). First, we increase group weights based on training loss. For each group $k$,
$w_{k}^{\prime}=w_{k}^{t-1}\times e^{lr\times C_{k}\times L_{\text{avg}k}}.$ (4)
@@
$L_{\text{avg}k}=\frac{\sum_{p\text{ in group }k}{l_{\text{dr}}(p)}}{N_{t}},$ (5)
where $N_{t}$ is the number of training data in step $t$. If all training data in a step belongs to a single group, $L_{\text{avg}k}$ for that group will be the average loss, and other group weights will remain unmodified.
Then, we normalize the modified weights and calculate the weight at step $t$. For each group $k$,
@@
During training, we aim to increase the weights of groups with higher contrastive loss, which leads the model to draw more attention to groups that are difficult to learn or contain more information.
@@
SPAR AugTriever coCondenser Contriever $*{}^{*}$ WebDRO ${}^{\sharp}$ ${}^{\flat}$ WebDRO
MS MARCO 22.8 19.2 20.6 16.2 20.6 26.15 $*{}^{*}$ 36.47 ${}^{\sharp}$ ${}^{\sharp}\flat$
trec-covid 65.6 53.1 53.5 40.4 27.4 72.16 69.26 72.43 73.98 ${}^{\sharp}\flat$
nfcorpus 32.5 26.4 30.3 28.9 31.7 30.70 30.74 29.22 29.75 ${}^{\sharp}\flat$
fiqa 23.6 18.5 22.3 25.1 24.5 23.79 $*{}^{*}$ 26.72 ${}^{\sharp}$ ${}^{\sharp}$
arguana 31.5 42.0 39.1 44.4 37.9 28.39 $*{}^{*}$ 47.59 ${}^{\sharp}$ 48.02
touche2020 36.7 26.1 21.6 11.7 19.3 21.85 $*{}^{*}$ 23.16 24.06 ${}^{\sharp}\flat$
quora 78.9 70.4 82.7 82.1 83.5 85.57 83.36 85.28 85.48 ${}^{\sharp}$
scidocs 15.8 13.4 14.7 13.6 14.9 13.42 $*{}^{*}$ 14.95 15.41 15.30
scifact 66.5 62.6 64.4 56.1 64.9 58.84 $*{}^{*}$ 57.54 ${}^{\sharp}$ ${}^{\sharp}$
nq 32.9 26.2 27.2 17.8 25.4 28.83 $*{}^{*}$ 42.48 ${}^{\sharp}$ ${}^{\sharp}\flat$
hotpotqa 60.3 57.2 47.9 34.0 48.1 53.81 52.27 50.25 ${}^{\sharp}$ ${}^{\sharp}$
dbpedia-entity 31.3 28.1 29.0 21.5 29.2 34.61 $*{}^{*}$ 34.01 ${}^{\sharp}$ ${}^{\sharp}\flat$
fever 75.3 56.9 59.7 61.5 68.2 72.15 $*{}^{*}$ 68.42 ${}^{\sharp}$ 70.93
climate-fever 21.4 16.4 17.7 16.9 15.5 18.87 21.05 18.45 18.95 18.89
cqadupstack 29.9 27.9 27.1 30.9 28.4 28.82 $*{}^{*}$ 33.64 ${}^{\sharp}$ ${}^{\sharp}$
trec-news 39.8 - - 25.4 35.2 41.18 40.24 39.68 ${}^{\sharp}$ ${}^{\sharp}$
signal1m 33.0 - - 21.4 25.0 22.99 $*{}^{*}$ 26.48 27.56 27.82
bioasq 46.5 - - 22.7 32.7 33.88 $*{}^{*}$ 30.07 33.73 ${}^{\sharp}\flat$
robust04 40.8 - - 29.8 32.7 41.01 37.60 ${}^{\flat}$ 43.23 ${}^{\flat}$
Avg. on BEIR14111Average score of 14 public datasets in BEIR. 43.0 36.2 37.0 33.4 36.0 40.84 42.15 43.15 45.14 46.03
Avg. All 41.3 - - 31.6 35.0 38.80 39.91 41.11 43.09 43.89
Table 1: Retrieval results of different models on MS MARCO and BEIR. All models are trained in the unsupervised setting and are evaluated using nDCG@10. The best result on each task is marked in bold, and the best result among unsupervised dense retrievers is underlined222 * indicate statistically significant improvements over Anchor-DR in unsupervised setting. $\sharp$ and $\flat$ indicate that over T5-base and Anchor-DR respectively in finetuned setting..

### v4

# Enhancing Dense Retrievers’ Robustness with Group-level Reweighting

@@ Abstract
The anchor-document data derived from web graphs offers a wealth of paired information for training dense retrieval models in an unsupervised manner. However, unsupervised data contains diverse patterns across the web graph and often exhibits significant imbalance, leading to suboptimal performance in underrepresented or difficult groups. In this paper, we introduce WebDRO, an efficient approach for clustering the web graph data and optimizing group weights to enhance the robustness of dense retrieval models. Initially, we build an embedding model for clustering anchor-document pairs. Specifically, we contrastively train the embedding model for link prediction, which guides the embedding model in capturing the document features behind the web graph links. Subsequently, we employ the group distributional robust optimization to recalibrate the weights across different clusters of anchor-document pairs during training retrieval models. During training, we direct the model to assign higher weights to clusters with higher loss and focus more on worst-case scenarios. This approach ensures that the model has strong generalization ability on all data patterns. Our experiments on MS MARCO and BEIR demonstrate that our method can effectively improve retrieval performance in unsupervised training and finetuning settings. Further analysis confirms the stability and validity of group weights learned by WebDRO. All codes will be released via GitHub.
Enhancing Dense Retrievers’ Robustness with Group-level Reweighting
@@ 1 Introduction
Dense retrieval models, which encode the semantics of text into dense vectors, have become prevalent and achieved state-of-the-art performance on many public benchmarks thanks to the surge of pre-trained language models like BERT Devlin et al. (2019) and T5 Raffel et al. (2020). With enough annotated data, dense retrieval models can adeptly capture semantic features in supervised scenarios, outperforming traditional sparse methods such as BM25 Robertson et al. (2009, 1995), which solely retrieves documents containing keywords already present in the query. However, annotated data is inaccessible in many scenarios either due to the extensive effort needed for annotation or privacy concerns. This highlights the necessity for unsupervised dense retrievers. Several paradigms of unsupervised dense retrieval have been proposed, among which leveraging web anchors is a promising choice. Primarily, anchor-document pairs contain weak supervision signal that indicates overlapping information Xie et al. (2023a); Zhang et al. (2020) and thus can approximate real search queries better than plain-text-based methods like ICT Izacard et al. (2021). In addition, anchor resources are abundant across the Internet, bearing greater potential for the scalability of retrievers compared to methods based on the title and abstract of the passage.
Although the web graph contains rich data patterns, the imbalanced data distribution is an inherent flaw of unsupervised retrieval data. Training robust dense retrievers with these data is challenging, as different patterns exhibit significantly different frequencies in training data, leading to imbalanced representation. Due to insufficient exposure to underrepresented data patterns, the model’s performance on downstream tasks involving similar patterns is hindered. To mitigate this, previous research has primarily focused on instance-level data selection to identify data points whose patterns haven’t been sufficiently trained by the model. To reach a more tailored selection, these approaches typically leverage feedback signals during the training process to assign a weight to each instance, such as Active Learning Settles (2011); Coleman et al. (2020) and Meta Adaptive Ranking Sun et al. (2021). These methods require the model to be trained multiple times, which is highly time-consuming and computationally intensive.
We address the lack of both efficient and effective retrieval data reweighting methods by introducing a group-level clustering and reweighting strategyWebDRO. WebDRO enables us to conduct data reweighting in a single pass, significantly saving computational resources. Additionally, data clustered into groups encapsulates much more information than a single data point and represents the common characteristics of the same pattern, WebDRO can enhance the quality and stability of reweighting.
@@
Specifically, we employ a two-step approach to train an unsupervised dense retrieval model using WebDRO. The first step (which we denote as Step A) is to generate clusters from unsupervised web data. We utilize structural information from the web graph to train an embedding model with contrastive learning (Step A-1). By distinguishing related web pages from unrelated ones, the embedding model can work as a link predictor that captures the semantic features of web documents. Leveraging the embedding model, we generate vector representations for documents in the anchor-document pairs and produce high-quality clusters through K-Means (Step A-2). In the second step (which we denote as Step B), we utilize Group Distributionally Robust Optimization (GroupDROOren et al. (2019); Sagawa et al. (2019) to reweight the clusters during the contrastive training of the retrieval model. Originally designed to tackle distributional shifts in language model training, GroupDRO dynamically adjusts the weights of different groups based on the training loss. It increases the weight of the groups with large training loss, therefore highlighting the underrepresented groups that the model has difficulty learning from and enhancing the model’s robustness in versatile retrieval tasks.
@@
We conduct experiments on two widely used retrieval datasets, MS MARCO Nguyen et al. (2016) and BEIR Thakur et al. (2021). Results indicate that our method notably enhances retrievers’ performance in both unsupervised and finetuned settings. Specifically, WebDRO yields a performance improvement of 1.2% (measured by nDCG@10) on MS MARCO and 1.1% on BEIR compared to our main baseline, Anchor-DR Xie et al. (2023b). Analytical experiments further demonstrate that our model consistently produces stable, high-quality clusters and accurately identifies groups that baseline dense retrievers struggle with.
The main contributions of our work are:
  • We explored web anchors’ usage in training dense retrievers and proposed WebDRO, a clustering and group-level reweighting algorithm for unsupervised dense retriever training.
  • Empirical analysis shows that WebDRO can efficiently enhance dense retrievers’ robustness, improving their performances on both unsupervised and finetuned settings.
  • We further confirmed WebDRO’s ability to efficiently identify underrepresented and challenging groups, thereby helping the model improve its capabilities more effectively.
@@ 2 Related Work
This section describes relevant work, focusing on dense retrieval models and distributionally robust optimization.
Dense Retrieval. Dense Retrieval is an approach where both the query and document are separately encoded through a neural network. The prevalent training approach, introduced by Karpukhin et al. (2020), trains the network with a contrastive loss. Although dense retrieval has outperformed sparse methods Robertson et al. (2009) in many tasks, its benefits are generally confined to well-annotated data. Unsupervised dense retrieval efforts focus on using unannotated data to generate query-document pairs that semantically resemble contrastive pairs in supervised scenarios. For instance, ICT Izacard et al. (2021) samples a text span in a passage as the query and the rest part as the document, and Co-docGao and Callan (2022) independently sample two text spans as the query and document. More recently, researchers have exploited unsupervised data at a greater scale and with more diverse constructing methods, namely E5 Wang et al. (2022) and BGE M3 Chen et al. (2024). In our work, we focus on one specific method of constructing contrastive pairs. Following Anchor-DR Xie et al. (2023b), we use anchors and corresponding webpages in the Internet to train dense retrievers. Web anchors are proved to be a high-quality data source for unsupervised retrievers, and has great potential of scaling up, considering the ubiquity and large number of anchors available in the Internet.
@@
Figure 1: Illustration of key steps in WebDRO. Figure 1(a) shows the clustering process in Step A, where we use web links to train an embedding model and use the model to cluster documents. Figure 1(b) shows the training process of the retrieval model in Step B. We aim to align the anchor with the corresponding document while pushing apart unrelated ones. Each group is assigned a weight that is dynamically updated based on the training loss.
@@ 3 Methodology
In this section, we first describe the preliminary of dense retrieval (Sec. 3.1). Then we introduce our WebDRO method, which is shown in Figure 1. In Step A, we use the method in Sec. 3.2 to train an embedding model using the web links. Then, we cluster the anchor-document pairs based on vector representations produced by the embedding model. In Step B, we train a robust dense retrieval model with Group Distributionally Robust optimization (GroupDRO) as introduced in Sec. 3.3.
@@ 3.1 Preliminary of Dense Retrieval
In dense retrieval, we train one single model to encode both queries and documents as dense vectors and then conduct relevance modeling in an embedding space. We use cosine similarity to measure the relevance between queries and documents. Through contrastive learning Oord et al. (2018); Chen et al. (2020); He et al. (2020), the encoder is optimized to align the queries with related documents and pull away the irrelevant documents from queries. Specifically, the following loss function is minimized during training:
$l_{\text{dr}}(q)=-\log\frac{e^{sim(E(q),E(p^{+}))}}{e^{sim(E(q),E(p^{+}))}+\sum_{i=1}^{k}{e^{sim(E(q),E(p^{-}_{i}))}}},$
@@
3.2 Step A: Clustering with the Embedding Model
@@
This subsection introduces our web graph based anchor-document cluster method. It learns the document representations using the document links in the web graph and then clusters the documents according to their semantic features.
@@
Document Relations in Web Graph. The web graph models the document relationship with hyperlinks, which allows users to navigate from one web page to another or within the same web page. It is evident that the documents linked with a hyperlink will likely share similar or related information. As shown in Figure 1(a), these four documents are related and are linked with the anchor hyperlink, e.g. the link between document A and document B indicate that “horses and gorillas have some similarities”, showing an analogy relationship; the links between document B and document C indicate the definitions of “horses” and “herbivorous”, showing the hypernymy relationship.
@@
Embedding Model Training. We train an embedding model using the document relations in the web graph. We encode the documents using a pretrained language model encoder and contrastively train the encoder. A web page can contain multiple links, so we enumerate each link to construct contrastive pairs as training examples. For each link, we regard the source web page as the query, the target web page as the positive document, and sample unlinked documents as negative documents. The objective function is identical to that in Sec. 3.1. As a result, the embedding model is trained as a link predictor that can distinguish linked web pages from non-linked ones.
@@
Document Clustering. Since the embedding model is capable of capturing webpages’ features, we use it to encode all documents as dense vectors. Then we use Mini-Batch K-Means Sculley (2010), a variant of K-Means Krishna and Murty (1999) to cluster embedding vectors of documents into a fixed number of groups. Some documents have more than one anchor that links to it, these anchor-document pairs are naturally clustered into the same group. To make sure each group has statistical significance, we set a threshold of the number of documents in a group as $MinSize=128$. All the groups with sizes smaller than $MinSize$ will be aggregated together and considered as a larger group. As this large group contains heterogeneous elements with different topics, we should not assign a single weight to represent this group. Therefore, the large group will not be considered in the reweighting discussed in Sec. 3.3.
3.3 Step B: Robust Dense Retriever Training
For each anchor in our corpus, we regard the text of the anchor as the query, and the content of the linked webpage as positive document. The anchor-document pairs contain different types of documents, and the distribution of different types could vary greatly: some occur frequently, while others are rare; some are easy for the model to learn, while others are challenging. Treating all data equally may affect the performance of dense retrievers when faced with underrepresented data types. To alleviate this problem, we follow Sagawa et al. (2019) to reweight different anchor-document groups and train more robust dense retrievers.
@@ 3.3 Step B: Robust Dense Retriever Training
Following GroupDRO, each group is assigned with a weight that indicates its importance, and the weights are dynamically updated during the training process. Specifically, we denote $n$ as the number of groups after clustering. Then, for a training instance $p$ in step $t$ that belongs to group $k$, the training loss of WebDRO is calculated as:
$L_{\text{WebDRO}}(p)=l_{\text{dr}}(p)\cdot w_{k}^{t}\cdot n\cdot C_{k},$ (2)
where $l_{\text{dr}}$ is the contrastive loss calculated using Eq. 1, $w_{k}^{t}$ is the weight of the $k$-th group in step $t$, and $C_{k}$ is size factor of the $k$-th group. We will next describe $C_{k}$ and $w_{k}$ respectively.
@@
Size Factor ($C_{k}$). In a standard GroupDRO scenario, data are sampled from each group according to the uniform distribution. However, the size of different anchor-document groups in our method may vary significantly. To accommodate this circumstance, WebDRO introduces a size factor $C_{k}$:
$C_{k}=\frac{1}{N_{k}}\cdot\frac{\sum_{j=1}^{n}{N_{j}}}{n},$ (3)
@@
where $N_{k}$ is the size of the $k$-th anchor-document group. Smaller groups will have a larger $C_{k}$ value, which serves as a balance for different group sizes. This is a natural extension of the original DRO, as $C_{k}=1$ for all $k$s when all groups have the same size (all $N_{k}$s are the same).
@@
Group Weighting ($w_{k}$). The initial weight of all groups is the same: $w_{k}^{0}=\frac{1}{n}$. In each step, the groups’ weights are updated before the calculation of $L_{\text{WebDRO}}$ (Eq. 2). First, we increase group weights based on training loss and get intermediate weights $w_{k}^{\prime}$. For each group $k$,
$w_{k}^{\prime}=w_{k}^{t-1}\cdot e^{lr\cdot C_{k}\cdot L_{\text{avg}k}}.$ (4)
@@
$L_{\text{avg}k}=\frac{\sum_{p\text{ in group }k}{l_{\text{dr}}(p)}}{N^{t}},$ (5)
where $N^{t}$ is the number of training data in step $t$. This is a natural extension of the case when all training data in a step belongs to a single group. $L_{\text{avg}k}$ for that group will thus be the average loss, and other group weights will remain unmodified.
After each step, the sum of group weights should be $1$, so we normalize the intermediate weights to get $w_{k}^{t}$. For each group $k$,
@@
During training, WebDRO aims to increase the weights of groups with higher contrastive loss, which leads the model to draw more attention to groups that are challenging for it. This enhances the model’s robustness on underrepresented groups in the source, thereby implicitly improving the model’s generalization ability on unseen target queries. Algorithm 1 shows the process of weight updating, which is the same as the above description.
Algorithm 1 Weight Updating in WebDRO
1:Input: Model $M$, step number $t$, number of groups $n$, learning rate for weight updating $lr$, size factor $C[1,\dots,n]$, size of batch $N^{t}$, data batch $p[1,\dots,N^{t}]$, original weights $w^{t-1}[1,\dots,n]$
2:Output: Updated weights $w^{t}[1,\dots,n]$
3:
4:$L_{\text{avg}}[1,\dots,n]\leftarrow[0,\dots,0]$
5:for $i=1$ to $N^{t}$ do
6:     Find the group number $k$ of $p[i]$
7:     Calculate the contrastive loss $l_{dr}(M,p[i])$
8:     $L_{\text{avg}}[k]\leftarrow L_{\text{avg}}[k]+l_{dr}(M,p[i])$
9:end for
10:for $k=1$ to $n$ do
11:     $L_{\text{avg}}[k]\leftarrow L_{\text{avg}}[k]\ /\ N^{t}$
12:     $w^{\prime}[k]\leftarrow w^{t-1}[k]\cdot\exp(lr\cdot C[k]\cdot L_{\text{avg}}[k])$
13:end for
14:$w_{\text{sum}}\leftarrow\sum_{k=1}^{n}{w^{\prime}[k]}$
15:for $k=1$ to $n$ do
16:     $w^{t}[k]\leftarrow w^{\prime}[k]\ /\ w_{\text{sum}}$
17:end for
18:RETURN $w^{t}[1,\dots,n]$
@@
SPAR AugTriever coCondenser E5-Large Anchor-DR WebDRO T5-Base Anchor-DR WebDRO
# Training Pairs - 22.6M 52.4M 8.8M 270M 13.8M 13.8M - - -
MS MARCO 22.8 19.2 20.6 16.2 26.2 26.15 27.37 36.47 39.13 40.58
trec-covid 65.6 53.1 53.5 40.4 61.8 72.16 69.26 72.43 73.98 78.04
nfcorpus 32.5 26.4 30.3 28.9 33.7 30.70 30.74 29.22 29.75 31.24
fiqa 23.6 18.5 22.3 25.1 43.2 23.79 25.12 26.72 28.11 28.45
arguana 31.5 42.0 39.1 44.4 44.4 28.39 31.27 47.59 49.18 48.02
touche2020 36.7 26.1 21.6 11.7 19.8 21.85 24.05 23.16 24.06 27.56
quora 78.9 70.4 82.7 82.1 86.1 85.57 83.36 85.28 85.48 85.85
scidocs 15.8 13.4 14.7 13.6 21.8 13.42 14.93 14.95 15.41 15.30
scifact 66.5 62.6 64.4 56.1 72.3 58.84 61.88 57.54 61.06 62.16
nq 32.9 26.2 27.2 17.8 41.7 28.83 34.48 42.48 45.24 47.23
hotpotqa 60.3 57.2 47.9 34.0 52.2 53.81 52.27 50.25 56.90 57.39
dbpedia-entity 31.3 28.1 29.0 21.5 37.1 34.61 35.34 34.01 36.39 38.09
fever 75.3 56.9 59.7 61.5 68.6 72.15 73.42 68.42 71.69 70.93
climate-fever 21.4 16.4 17.7 16.9 15.7 18.87 21.05 18.45 18.95 18.89
cqadupstack 29.9 27.9 27.1 30.9 38.9 28.82 32.97 33.64 35.79 35.25
trec-news 39.8 - - 25.4 - 41.18 40.24 39.68 43.05 41.80
signal1m 33.0 - - 21.4 - 22.99 24.36 26.48 27.56 27.82
bioasq 46.5 - - 22.7 - 33.88 38.45 30.07 33.73 35.86
robust04 40.8 - - 29.8 - 41.01 37.60 44.19 43.23 43.51
Avg. on BEIR14111Average score of 14 public datasets in BEIR. 43.0 36.2 37.0 33.4 44.2 40.84 42.15 43.15 45.14 46.03
Avg. All 41.3 - - 31.6 - 38.80 39.91 41.11 43.09 43.89
Table 1: Retrieval results of different models on MS MARCO and BEIR. All models are evaluated by nDCG@10. Model names in the "Finetuned Dense Retrievers" category only indicate the base model before finetuning. The best result on each task is marked in bold, and the best result among unsupervised dense retrievers is underlined.